Fermi National Laboratory


Accelerator Update

Friday August 29
The day shift began with Tevatron (TeV) experts conducting studies, with the Antiprotons Source (Pbar) stacking, and with MiniBooNE taking beam.

Recycler experts took one shot of beam to adjust their BPM timing at 9:21 AM.

By 11:36 AM, Operations had received reports of small LCW leaks from the Main Injector (MI), TeV, and Debuncher. TeV studies ended at 11:37 AM and TeV technicians started their LCW repairs.

Operations put the TeV through a dry squeeze at 12:51 PM.

At 1 PM, a Linac quadrupole power supply began tripping off. Experts changed out a control card and had the supply working twenty minutes later.

Operations put the TeV into shot setup at 2:19 PM.

A Pbar water expert began his repairs at 3:57 PM.

At 4:01 PM, the TeV suffered a sector D1 quench. Experts investigated the possible causes for the rest of the day and evening shift.

Operations prepared the Pbar Prevault area for access at 4:18 PM. The Pbar experts were out of access thirty minutes later.

A Pbar RF station (ARF4) had developed a problem at the very end of the last shot. An expert repaired the station's driver at 4:19 PM.

Pbar resumed stacking at 5:06 PM.

A MI and Switchyard expert began conducting a 120 GeV slow extraction study at 6:06 PM. The studies ended around midnight.

A NuMI expert began a study period at 6:07 PM. The studies ended around midnight.

Operation loaded a 12x0 TeV study-store at 11:44 PM.

Saturday August 30
The midnight shift began with TeV experts continuing their investigation of the quench, with Pbar stacking, and with MiniBooNE taking beam.

Operations loaded at 36x0 TeV study-store 1:29 AM.

At 2 AM, Booster beam was not making it into MI. Initial investigations indicated at kicker problem. However, while investigating, operators unplugged a scope cable and things began to work again. They continued to work after the cable was plugged back in.

The TeV experts completed their studies at 4:51 AM.

Operations put the TeV through a dry squeeze at 5:04 AM.

Operations put the TeV into shot setup around 5:30 AM. Unfortunately, the TeV suffered at sector A1 quench at 6:52 AM. Pbar resumed stacking.

At 10:35 AM, two Booster RF stations (BRF13 & 14) tripped off due to bad vacuum. These stations reset okay.

Both MI and Pbar studies were completed by 12:18 PM.

Operations loaded another 12x0 study-store to help TeV experts investigate our quench problems. At 4:24 PM, Operations loaded 36x4 study-store. Operations terminated the study-store at 7:30 PM. TeV experts reported that the quenches were due to a problem with cogging and with the low level RF.

Two Booster RF stations (BRF 15 & 16) tripped off due to a vacuum pump tripping off. The pump and stations all reset okay.

At 10:46 PM, MI BLT problems occurred. Operations called in experts to investigate. They adjusted some possible timing conflicts at 11:49 PM.

Operations loaded another 36x4 study-store 10:51 PM.

Operations discovered a Pbar scope at 11:46 PM that had died. Experts were called in to replace the scope.

Sunday August 31
The midnight shift began with TeV expert investigating their quench problems, with Pbar stacking, and with MiniBooNE taking beam.

At 1:04 AM, the experts working on the MI BLT reported that they updated its communication capabilities.

Operations terminated the TeV study-store at 1:25 AM.

Operations put the TeV into shot setup and established store 2975 at 4:43 AM. Its initial luminosity was 2.8E31. Pbar resumed stacking.

A Pbar Debuncher RF station (DRF1) tripped off at 5:09 AM. Since this station being off wasn't hindering stacking, Operations decided to leave it off for the time being.

At 7:18 AM, a MI RF station (MIRF16) had already tripped off several times due to LCW flow problems. The station reset each time. Operators think it may have been due to a bad flow meter.

At 11:55 AM, the MI beam was halted due to a MECAR problem that occurred when NuMI experts began a study period. Operators power cycled the device locally and this had the problem resolved thirty minutes later. Pbar resumed stacking.

A Linac RF station (LRF4) tripped off at 1:47 PM. It reset okay.

NuMI studies finished at 6:42 PM.

LRF4 tripped off 6:47 PM and again at 7:11 PM. It reset okay.

Operations terminated store 2975 at 9:06 PM.

Operators made the D-zero (D0) collision hall ready for access at 9:15 PM.

Operations shot some protons into the Recycler for studies at 10:23 PM.

D0 came out of access at 11:04 PM.

Operations put the TeV through a dry squeeze at 11:38 PM.

Monday September 1
The midnight shift began with the TeV in a dry squeeze, with Pbar stacking, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 12:10 AM.

At 1:32 AM, Operation reported still getting MI BLT errors. Experts were notified.

Operations established store 2977 at 2:03 AM with an initial luminosity of 4.2E31.

Today's Plan
The plan for Monday was to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/2/2003   email Fermilab